Determine the vertical change of a line with an x-intercept at (-4, 0) and a y-intercept at (5, 0).

Found 2 solutions by josmiceli, Theo:
Answer by josmiceli(19441)

The problem has an error, maybe in copying it.
(5,0) is not a y-intercept
The y-intercept has to look like: ( 0, y )
With any real value for +y+

Answer by Theo(13342)

the line starts off at (-4,0) and ends at (5,0)
the slope is equal to (y2-y1) / (x2-x1)
assign one of the points to (x1,y1) and assign the other point to (x2,y2) and plug those values into the formula and solve for the slope.
You can assign any point to (x1,y1) and any other point to (x2,y2). the only requirement is that both points be on the same line.
I did the following:
(x1,y1) = (-4,0)
(x2,y2) = (5,0)
PROBLEM !!!!!
(5,0) is not a y-intercept !!!!!
If that point is a y-intercept, it has to be (0,5).
the y intercept is the value of y when the value of x is equal to 0.
the x intercept is the value of x when the value of y is equal to 0.
looks like you have 2 x intercepts here.
i changed (5,0) to (0,5) in order to show you how to do it.
your corrected points are:
(x1,y1) = (-4,0)
(x2,y2) = (0,5)
y2-y1 = 5 - 0 = 5
x2-x1 = 0 - (-4) = 0 + 4 = 4
your slope is equal to 5/4.
to create the equation of the line connecting the 2 points, you need to find the y intercept.
one way of doing that is to replace x with the value of the x coordinate of one of the points and replace y with the value of the y coordinate of that point and then solve for b in the equation of:
y = mx + b
m is the slope = 5/4
replace x with 0 and y with 5 and you get:
5 = 4/5*0 + b
solve for b to get:
b = 5
the equation of your line is:
y = 5/4x + 5
here's the graph of your line.
you can see that the points (-4,0) and (0,5) are both on the line.
$$$